Use standard reduction potentials to calculate the equilibrium constant for the reaction: Cd2+(aq) + 2Fe2+(aq)Cd(s) + 2Fe3+(aq) Hint: Carry at least 5 significant figures during intermediate calculations to avoid round off error when taking the antilogarithm. Equilibrium constant: _____G° for this reaction would be(greater/less) than zero
Answer #1

E° = E°c - E°a

= -0.403 - (0.771)

= -1.174 V

-1.174 =( 0.0591/2)× log(K)

K = 1.87×10-40

∆G° is less greater than one